The lenght of a rectangle is 5m more than it's breadth. find the lenght and the breadth If perimeter is 70 m
1
Expert's answer
2013-06-26T09:59:23-0400
Perimeter=2* (length + breadth)
X – breadth
X + 5 – length
Perimeter = 2 * ( X +5 + X ) = 70
2 * ( 2X + 5 ) = 70
2X + 5 =70/2
2X + 5 = 35
2X = 35 – 5
2X= 30
X = 15
15 m - breadth
15 + 5 = 20 m – length
Answer: 15 m – breadth , 20 m – length .